can someone please help me with this problem:
Evaluate the function at the given value of the independent variable and simplify.

f(x) = x^3 + 3 / x^2 - 3 ; f(3)

Answer by checkley77(12569)

f(x) = x^3 + 3 / x^2 - 3 ; f(3)
f(3)=3^3+3/3^2-3
f(3)=27+3/9-3
f(3)=27+1/3-3
f(3)=(27*3+1+3*3)/3
f(3)=(81+1+9)/3
f(3)=91/3 answer.
OR:
f(3)=3^3+3/(3^2-3)
f(3)=27+3/(9-3)
f(3)=27+3/6
f(3)=(27*6+3)/6
f(3)=(162+3)/6
f(3)=165/6
f(3)=55/2 answer.